OpenStudy (anonymous):

-8x+9x-21=-12-16 What is the solution to the equation?

OpenStudy (anonymous):

-8x+9x-21=-12-16 x-21=-12-16 x-21=-28 +21 +21 x=-7

OpenStudy (anonymous):

thanks again im just not good at this stuff